# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v38

@@ 6.2 Symmetric Equilibrium
There exists $\alpha^{*}>0$ such that $\alpha^{*}{\bf Q}$ is an equilibrium of the canonical game. Equivalently, $\alpha^{*}{\bf L}^{-1}{\bf Q}$ specifies an equilibrium of the original trading game, in the canonical form of Corollary 6.3.
Example 6.6.
(Theorem 6.5, Binary Signal Case)
Suppose the signal is binary, $S=\{s_{1},s_{2}\}$. Substituting the equilibrium postulate (31) $D^{*}=\frac{1}{2}\begin{bmatrix}\alpha&-\alpha\\ -\alpha&\alpha\end{bmatrix}$ into the insider’s first-order condition (30) conditional on, say, $s_{1}$ gives
$\begin{bmatrix}1\\ 0\end{bmatrix}-\begin{bmatrix}\mathbb{E}[p_{1}]\\ 1-\mathbb{E}[p_{1}]\end{bmatrix}-\alpha^{2}\begin{bmatrix}\mathbb{E}[p_{1}p_{2}]\\ -\mathbb{E}[p_{1}p_{2}]\end{bmatrix}=(1-\mathbb{E}[p_{1}]-\alpha^{2}\mathbb{E}[p_{1}p_{2}])\begin{bmatrix}1\\ -1\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$ (34)
where $\mathbb{E}[\,\cdot\,]$ is the expectation taken under $D^{*}$, and $p_{k}$ is the posterior probability of $s_{k}$, $k=1,2$. Substituting $D^{*}$ into the market maker’s posterior (26) gives
@@
$(p_{1},p_{2})\stackrel{{\scriptstyle d}}{{\sim}}(\frac{e^{Z}}{e^{Z}+1},\frac{1}{e^{Z}+1}),\quad\mbox{ where }Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(\frac{\alpha^{2}}{2},2\alpha^{2}).$ (35)
By (35), the quantities $\mathbb{E}[p_{1}]$ and $\mathbb{E}[p_{1}p_{2}]$ in (34) are moments of a logit-normal distribution and functions of $\alpha$. Write $\mathbb{E}[p_{1}]$ as $\phi_{1}(\alpha)$ and $\mathbb{E}[p_{1}p_{2}]$ as $\phi_{2}(\alpha)$. Define $\Phi(\alpha)=1-\phi_{1}(\alpha)-\alpha^{2}\phi_{2}(\alpha)$. Then the first-order conditions (34) conditional on $s_{1}$ and $s_{2}$ become, respectively,
[…]
Collating these two symmetric equations side-by-side yields the matrix equation $\Phi(\alpha){\bf Q}=0$ of (33) when $K=2$. To find equilibrium, it suffices to find $\alpha^{*}>0$ such that $\Phi(\alpha^{*})=0$. Since $\Phi(0)=\frac{1}{2}$ and $\lim\limits_{\alpha\rightarrow\infty}\Phi(\alpha)\uparrow 0$, we have $\Phi(\alpha^{*})=0$ for some $\alpha^{*}>0$. [^19] That $\lim\limits_{\alpha\rightarrow\infty}\Phi(\alpha)\uparrow 0$ follows from Fatou’s Lemma. The existence of $\alpha^{*}>0$ such that $\Phi(\alpha^{*})=0$ then follows directly from the Intermediate Value Theorem. This proves Theorem 6.5 when the signal is binary.
Intuition of $\alpha^{*}$
The endogenous constant $\alpha^{*}$ characterizing equilibrium is the insider’s optimal portfolio size in the canonical pseudo-trading game. The intuition behind $\alpha^{*}$ builds on the simplified setting discussed in Section 2—see again Figure 1b. This is demonstrated by Example 6.7.
Example 6.7.
In the same setting as Example 6.6, consider the insider’s utility and the market maker’s posterior (35), as $\alpha$ varies.
  • When $\alpha=0$, the insider has zero demand, yielding zero utility and revealing no information. The market maker’s posterior reduces to the prior. This cannot be an equilibrium—the insider’s marginal utility is strictly positive.
  • When $\alpha\rightarrow\infty$, the insider scales up his demand without bound, fully revealing his information and resulting in zero profit. The market maker’s posterior concentrates on the insider’s signal. This cannot be an equilibrium—the insider’s marginal utility is strictly negative.
  • When $\alpha=\alpha^{*}$, the insider’s marginal utility is zero, the market maker’s belief is correct, and equilibrium is obtained. The equilibrium posterior is biased towards the insider’s signal, partially revealing his information because the market maker correctly anticipates his trades.
Comparative Statics of $\alpha^{*}$
@@
$\alpha^{*}$ is determined only by the number of signals KK—it is independent of $\eta(\,\cdot\,|\,\cdot\,)$ or $(\sigma_{i})$. $\alpha^{*}(K)$ is an increasing concave function of $K$, as shown in Figure 2. Increasing the number of signals enhances the insider’s informational advantage over the market maker, resulting in larger insider trades. However, larger insider trades also introduce greater variation in informed demand across signals, diminishing the incremental gain in the insider’s information advantage.
@@
$S=\{s_{1},\ldots,s_{K}\}$ Set of possible private signal realizations.
$X=\{x_{1},\ldots,x_{N}\}$ Set of possible future states of the world (may be infinite).
$k,l$ Indices for signals/information types (e.g., volatility, tail risk, skewness).
$i,j$ Indices for states/AD securities (equivalently, option strikes via Breeden-Litzenberger).
$\eta(x_{i}\mid s_{k})$ Probability of state $x_{i}$ conditional on signal $s_{k}$ (signal-implied AD payoff distribution).
@@
$W^{*}(x_{i}\mid s_{k})$ Insider’s demand in security $i$ under signal $s_{k}$ (“load on these states, short those”).
$\varepsilon_{i}$ Noise trade in security $i$.
$\omega=(\omega_{i})_{i=1}^{N}$ Aggregate order-flow vector observed by the market maker, where $\omega_{i}=W^{*}(x_{i}\mid s_{k})+\varepsilon_{i}$.
$\sigma_{i}$ Noise-trade intensity in market $i$ (larger $\sigma_{i}$ implies noisier order flow).
@@
$\pi^{*}(s_{k}\mid\omega)$ Market maker’s posterior probability of signal $s_{k}$ after observing $\omega$.
$P^{*}(x_{i}\mid\omega)$ Equilibrium price of AD state claim $i$ conditional on $\omega$ (state price/pricing kernel).
$\Lambda=(\Lambda_{i,j})$ Equilibrium price impact matrix: contemporaneous sensitivity of $P^{*}(x_{i}\mid\omega)$ to order flow in security $j$; diagonals are within-market impact, off-diagonals are cross impact.
@@
Q Signal-tilt matrix: maps each signal into a long-short trade over states (see (32)).
L Information intensity matrix: summarizes how informative joint order flow is across contracts (see Definition 3).
${\bf L}^{-1}$ Liquidity adjustment matrix: converts signal tilts into equilibrium trading weights by adjusting for relative noise/liquidity.
$\alpha^{*}$ Equilibrium overall trading scale, capturing the profit–revelation trade-off (see Example 6.7).
$\beta^{(k)}$ Portfolio weights under signal $s_{k}$ (the $k$th column of $\alpha^{*}{\bf L}^{-1}{\bf Q}$), entering informed demand $W^{*}(\,\cdot\,\mid s_{k})$ (see (37)).
@@ 7 Price Discovery
For ease of reference, Table 1 collects the key equilibrium quantities, organized by the model’s information channel:
@@
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. Applying the Breeden–Litzenberger formula (7) expresses these Arrow–Debreu quantities in terms of options.
@@ 7.1 Informed Demand
  • Step 2 (Liquidity adjustment). Rotate and scale the signal tilt by the matrix ${\bf L}^{-1}$. Economically, $\bf L$} summarizes how informative order flow is to market makers when they observe joint trading across contracts; applying ${\bf L}^{-1}$ therefore damps positions in directions where order flow is more revealing (and hence more costly in terms of price impact).
@@
Recall that this same direction–aggressiveness decomposition already appears in the single-asset informed demand (5) of Kyle (1985). There, the insider decides whether to buy or sell based on whether his signal exceeds the prior mean, and then chooses how aggressively to scale that position given the resulting price impact. To make this parallel as transparent as possible, compare the canonical AD representation (Corollary 6.3) to the single-asset rule (5):
@@
Returning to our motivation, the portfolio characterization of informed demand provides an equilibrium foundation for a broad class of option strategies used in practice. To illustrate, we consider a few prominent strategies for trading on the level, volatility, and skewness of an underlying payoff. Market participants routinely express such views using vertical spreads, straddles, and skew trades, respectively. In contrast, existing theoretical treatments of option trading are largely confined to single-option environments, and broader multi-security frameworks either shut down strategic informed portfolio choice, or else rely on restrictive distributional structure; consequently, they cannot endogenously generate these nonlinear, multi-leg option portfolios observed in practice. [^21] SeeBack (1993a); Biais and Hillion (1994); Easley et al. (1998a); Collin-Dufresne et al. (2021); Admati (1985); Malamud (2015); Chabakauri et al. (2022).
@@
Relative to the single-asset case in Kyle (1985), where informed demand is linear, access to options reshapes informed demand into a nonlinear, multi-leg portfolio. Equilibrium demand tilts payoffs toward strikes where the two signal-implied distributions differ most and avoids unnecessary tail exposure.
@@
Portfolio intuition. Under the high-volatility signal $s_{1}$, the distribution places relatively more weight on tail outcomes. Accordingly, the informed demand loads on the wings and avoids the center (panel 4b), producing a straddle-type position—a wing-loaded long-volatility portfolio.
Conversely, under the low-volatility signal $s_{2}$ the payoff is more concentrated near the mean. Panel 4d shows that informed demand flips sign across strikes, producing a butterfly position: it concentrates exposure near the center and shorts wing payoffs that are unlikely to realize under low volatility.
Relative to single-security environments, volatility information is inherently cross-state and therefore multi-dimensional. Equilibrium correspondingly takes the form of a multi-option portfolio: volatility information is expressed as a center-versus-wings trade, yielding straddle- and butterfly-type positions as equilibrium outcomes.
@@
Skew trades are a core building block in option markets, commonly implemented via ratio spreads and closely related strategies. [^24] See Hull (2015); Natenberg (2014); McMillan (2012); Vine (2011); Overby (2007). Option-implied skewness and skew premia are likewise central in the empirical literature. [^25] See Dennis and Mayhew (2002); Bakshi et al. (2003); Conrad et al. (2013b); Bollen and Whaley (2004); Kozhan et al. (2013).
Beyond underscoring the practical scope of our model, the above examples also sharpen the point that derivatives are central to price discovery. When private information departs from the mean, it corresponds to a cross-state/strike payoff tilt that is inherently high-dimensional and, generically, cannot be spanned by a single contract; cf. Proposition 3.3(ii). Derivatives matter precisely because they span—and therefore can reveal—such tilts. Accordingly, a general specification such as ours is both natural and necessary: it allows equilibrium to translate arbitrary information into coherent portfolios across strikes.
@@ 7.2 Price Impact
We next characterize the equilibrium price impact matrix $\Lambda$. Its $(i,j)$ element $\Lambda_{i,j}$ is the equilibrium marginal effect of trades in $x_{j}$ on the price of $x_{i}$ (diagonals are within-market impact; off-diagonals are cross impact). For the reader’s convenience, Table 1 summarizes the notation used in this subsection.
@@
Here, $\eta(x_{i}\mid\,\cdot\,)$ is the signal-by-signal payoff profile of claim $x_{i}$. Given order flow $\omega$, the market maker forms a posterior over signals. Throughout the sequel, we write ${\operatorname{\mathbb{Cov}}}(\cdots,\cdots\mid\omega)$ for the (conditional) covariance under that posterior. The outer expectation $\mathbb{E}[\,\cdot\,]$ averages over order flow $\omega$ under its equilibrium distribution. Equivalently, the realized (order-flow-conditional) cross price impact is obtained by dropping the outer expectation from (38):
@@
  • Signal-tilt coefficient $\big(1-\frac{1}{K}\big)$: This is the loading on the realized signal in the informed demand (37) portfolio recipe—Step 1. When there are many signals ($K$ is large), the realized signal carries less weight relative to the cross-signal average, dampening price impact.
  • Equilibrium aggressiveness $\alpha^{*}$: This is the overall scale of informed demand (37)—Step 3. If the insider trades more aggressively in equilibrium, a given marginal order is more likely to be information-driven, so prices respond more.
  • Noise adjustment $\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}$: This term captures the standard idea that more noise trading makes order flow less informative. More noise in either market dilutes how much a trade in $x_{j}$ moves beliefs (and thus prices) about $x_{i}$. Indeed, we already saw this intuition in the toy three-state model of Section 2 (see footnote 8).
@@
    is the (posterior) payoff covariance between securities $x_{i}$ and $x_{j}$—and thus captures whether the market maker’s inference from order flow tends to move their prices in the same direction:
@@
    (Negative cross price impact.) If security $x_{i}$ pays off only under signal $s_{k_{i}}$ and $x_{j}$ only under a different signal $s_{k_{j}}\neq s_{k_{i}}$, then cross impact is negative: $\Lambda_{i,j}<0$. This is the equilibrium analog of Example 5.3(ii): buy pressure in $x_{i}$ makes the market maker infer that $x_{j}$ is less likely to pay off (and vice versa). More precisely, the sign of $\Lambda_{i,j}$ is determined by the (order-flow-conditional) posterior payoff covariance term in (39), which is negative in this case:

### v39

@@ 6.2 Symmetry and One-Parameter Equilibrium
The canonical game admits an equilibrium $D^{*}=\alpha^{*}\,{\bf Q}$ in which the insider trades the active information contract against the average of the others, with endogenous trading scale $\alpha^{*}$. Reversing the information-unit conversion then yields the corresponding equilibrium of the original trading game with canonical form loadings matrix (23), ${\bf L}^{-1}D^{*}=\alpha^{*}\,{\bf L}^{-1}{\bf Q}.$
Intuition for $\alpha^{*}$
The intuition underlying Theorem 6.3 is the same as that highlighted in the toy setting of Section 2, now shown to be robust in full Arrow-Debreu generality. Raising $\alpha$ scales up the long-short portfolio and increases the insider’s exposure to the true story, but it also makes aggregate order flow more informative, which erodes the insider’s informational edge. Equilibrium $\alpha^{*}$ obtains at the knife-edge where this tradeoff balances.
The next subsection demonstrates the especially transparent binary-signal case (which nests the toy setting of Section 2), where inference collapses to a single log-odds statistic.
6.3 Example - Binary Signal Case
Suppose there are two possible information stories, $S=\{s_{1},s_{2}\}$. In the canonical game there is one contract per story, paying $1$ if that story is true and 0 otherwise. The natural informed trade is a market-neutral spread: under each signal the insider buys the “active” contract and shorts the alternative by the same amount. This long-short trade corresponds to the strategy matrix
$D^{*}=\frac{\alpha}{2}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix},\qquad\alpha>0,$
In this two-story case, the cross-section of information-contract prices $(p_{1},p_{2})$ is fully summarized by a single posterior log-odds statistic $Z$:
@@ 6.3 Example - Binary Signal Case
$(p_{1},p_{2})\stackrel{{\scriptstyle d}}{{=}}\left(\frac{e^{Z}}{e^{Z}+1},\,\frac{1}{e^{Z}+1}\right),\qquad\mbox{ where }Z\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}\!\left(\frac{\alpha^{2}}{2},\,2\alpha^{2}\right).$ (27)
Thus observing order flow is equivalent to observing the “evidence score” $Z$: large $Z$ is strong evidence for $s_{1}$, while large negative $Z$ favors $s_{2}$. [^18] See Lemma A.2 in the Appendix.
When the insider does not trade and order flow is pure noise ($\alpha=0$), the evidence score is degenerate at $Z=0$, so prices stay at the prior (no learning). When the insider scales the spread without bound ($\alpha\to\infty$), $Z$ becomes overwhelmingly informative and prices concentrate on the realized story, so informational rents vanish. More generally, trading scale $\alpha$ affects payoff only through the evidence score $Z$. Larger $\alpha$ scales up the spread, but it also makes ZZ—and hence order flow—more informative.
This tradeoff is characterized by the marginal indifference condition for $\alpha$ (our scalar equilibrium equation) [^19] See Equation (A.23) in the Appendix.
$\Phi(\alpha)\;\equiv\;1-\mathbb{E}[p_{1}]-\alpha^{2}\,\mathbb{E}[p_{1}p_{2}]\;=\;0,$ (28)
where the expectations are taken under (27) (and hence are functions of $\alpha$). Here $1-\mathbb{E}[p_{1}]$ is the average “wedge” the insider earns from the active story, while $p_{1}p_{2}=p_{1}(1-p_{1})$ is the posterior variance: it is largest when $Z\approx 0$ (the market is unsure) and small when $|Z|$ is large (the market is confident), making precise how marginal scaling of the spread further sharpens inference and erodes informational rents.
The equilibrium scale $\alpha^{*}$ solves $\Phi(\alpha)=0$, i.e., the marginal gain from increasing exposure is offset by the marginal information leakage. We have $\Phi(0)=\tfrac{1}{2}>0$, while $\Phi(\alpha)<0$ for sufficiently large $\alpha$, and moreover $\Phi(\alpha)\uparrow 0$ as $\alpha\to\infty$. [^20] The limit $\Phi(\alpha)\uparrow 0$ follows from Fatou’s Lemma. Existence of $\alpha^{*}>0$ with $\Phi(\alpha^{*})=0$ follows by the Intermediate Value Theorem. Hence there exists $\alpha^{*}>0$ solving $\Phi(\alpha)=0$, which proves Theorem 6.3 in this binary signal case.
6.4 Comparative Statics of $\alpha^{*}$
@@
We now return to the general model. The canonical-game reduction shows that the equilibrium trading scale $\alpha^{*}$ is pinned down solely by the number $K$ of signals and is therefore invariant to the particular signal-implied payoff distributions $\eta(\,\cdot\,|\,\cdot\,)$ and contract-level noise intensities $(\sigma_{i})$. Figure 2 plots the resulting mapping $K\mapsto\alpha^{*}(K)$.
As $K$ increases, there are more distinct payoff-distribution “stories” the market must disentangle from cross-contract order flow, so equilibrium informed trading becomes more aggressive. At the same time, a larger $K$ also increases how differently informed demand loads across stories, making order flow more diagnostic and steepening the informational (price-impact) cost of further scaling. The net effect is that $\alpha^{*}(K)$ rises with $K$ but with diminishing increments: $\alpha^{*}(K)$ is increasing and concave.
Put differently, as more distinct information stories are traded and priced through the cross section, they support larger informed positions. But each additional story contributes less to overall trading intensity, because the cross-sectional pattern of order flow becomes progressively more diagnostic of which story is active.
@@
$S=\{s_{1},\ldots,s_{K}\}$ Set of possible information stories (signals) about the payoff distribution (e.g., volatility, tail risk, skewness).
$X=\{x_{1},x_{2},\ldots\}$ Set of possible future states (e.g., terminal underlying prices); may be infinite.
$k,l$ Indices for information stories/signals.
$i,j$ Indices for Arrow-Debreu states/claims (e.g., option strikes in the options implementation via (7).)
$\eta(x_{i}\mid s_{k})$ Probability of state $x_{i}$ under story $s_{k}$ (i.e., the signal-implied distribution over states).
@@
$W^{*}(x_{i}\mid s_{k})$ Insider’s equilibrium order/position in claim $x_{i}$ under story $s_{k}$.
$\varepsilon_{i}$ Noise order in claim $x_{i}$ (non-informational order flow).
$\omega=(\omega_{i})_{i=1}^{N}$ Aggregate order flow across claims received by the market maker; $\omega_{i}=W^{*}(x_{i}\mid s_{k})+\varepsilon_{i}$.
$\sigma_{i}$ Noise-trade intensity in market $i$ (larger $\sigma_{i}$ means noisier, less informative order flow).
@@
$\pi^{*}(s_{k}\mid\omega)$ Market maker’s posterior probability that story $s_{k}$ is active, conditional on order flow $\omega$.
$P^{*}(x_{i}\mid\omega)$ Equilibrium state price (pricing kernel) for claim $x_{i}$ conditional on $\omega$.
$\Lambda=(\Lambda_{i,j})$ Equilibrium price-impact matrix: $\Lambda_{i,j}$ is the response of $P^{*}(x_{i}\mid\omega)$ to order flow in claim $x_{j}$ (own-impact on diagonals, cross-impact off-diagonals).
@@
Q Signal-tilt matrix: encodes the market-neutral long–short tilt that isolates the realized story (see (25)).
L Information intensity matrix: how diagnostic joint cross-contract order flow is about which story is active (see Definition 3).
${\bf L}^{-1}$ Information adjustment matrix: rotates/scales the raw signal tilt into equilibrium weights, downweighting directions where order flow is more revealing.
$\alpha^{*}$ Overall position scale, balancing profit vs. information revelation (see Section 6.3).
$\beta^{(k)}$ Story-$s_{k}$ equilibrium weights (the $k$-th column of $\alpha^{*}{\bf L}^{-1}{\bf Q}$) used to build $W^{*}(\,\cdot\,\mid s_{k})$ (see (30)).
@@ 7 Price Discovery
For ease of reference, Table 1 recollects the key equilibrium quantities, organized by our model’s information channel:
@@
The private signal $s_{k}$ induces equilibrium informed demand $W^{*}(\,\cdot\,\mid s_{k})$, which is part of aggregate order flow $\omega$. Conditioning on $\omega$, the market maker sets Arrow-Debreu state prices $P^{*}(\,\cdot\,\mid\omega)$; the within- and cross-price responses are collected in the price impact matrix $\Lambda$. Applying the standard formula (7) expresses these Arrow–Debreu quantities in terms of options.
@@ 7.1 Informed Demand
  • Step 2 (Liquidity adjustment). Rotate and scale the signal tilt by the matrix ${\bf L}^{-1}$. $\bf L$} summarizes how informative order flow is to market makers as they observe joint trading across contracts; applying ${\bf L}^{-1}$ therefore damps positions in directions where order flow is more revealing and hence more costly in terms of price impact.
@@
Recall that this same direction–aggressiveness decomposition already appears in the single-asset informed demand (5) of Kyle (1985). There, the insider decides whether to buy or sell based on whether his signal exceeds the prior mean, and then chooses how aggressively to scale that position given the resulting price impact. This parallel becomes very transparent when comparing the canonical form (as defined in (23)) $\alpha^{*}{\bf L}^{-1}{\bf Q}$ of our Arrow-Debreu equilibrium to the single-asset trading rule (5):
@@
Returning to our motivation, the portfolio characterization of informed demand provides an equilibrium foundation for a broad class of option strategies used in practice. To illustrate, we consider a few prominent strategies for trading on the level, volatility, and skewness of an underlying payoff. Market participants routinely express such views using options. In contrast, existing theoretical treatments of option trading are largely confined to single-option environments, or multi-security frameworks that either shut down strategic portfolio choice, or else rely on restrictive distributional assumptions; consequently, they cannot endogenously generate the nonlinear, multi-leg option portfolios observed in practice.
@@
Even when private information is only about the mean—the Kyle (1985) case—the availability of options reshapes informed demand into a nonlinear, multi-leg portfolio. Example 7.1 illustrates this.
@@
Volatility information concerns shifts in probability mass between the center and the tails, so equilibrium informed demand naturally takes a center–versus–wings form, yielding straddle- and butterfly-type positions. This is Example 7.2.
@@
Portfolio intuition. Under the high-volatility signal $s_{1}$, the distribution places relatively more weight on tail outcomes. Accordingly, the informed demand loads on the wings and avoids the center (panel 4b), producing a straddle-type position—a wing-loaded long-volatility portfolio. Conversely, under the low-volatility signal $s_{2}$ the payoff is more concentrated near the mean. Panel 4d shows that informed demand flips sign across strikes, producing a butterfly position: it concentrates exposure near the center and shorts wing payoffs that are unlikely to realize under low volatility.
@@
Skew is routinely traded in option markets—through ratio spreads and closely related structures. [^24] See Hull (2015); Natenberg (2014); McMillan (2012); Vine (2011); Overby (2007). Option-implied skewness and skew premia are likewise central in the empirical literature. [^25] See Dennis and Mayhew (2002); Bakshi et al. (2003); Conrad et al. (2013); Bollen and Whaley (2004); Kozhan et al. (2013). Example 7.3 shows that a pure skewness signal endogenously recovers this familiar skew trade.
@@
Beyond underscoring the practical scope of our model, the above examples also reinforce the point that derivatives are central to price discovery. When private information departs from the mean, it corresponds to a cross-strike payoff tilt that cannot be spanned by a single contract; cf. Proposition 3.3(ii). Derivatives matter precisely because they span—and therefore can reveal—such tilts. The breadth of observed option strategies points to information that is rarely one-dimensional; mean-only restrictions are mainly methodological. Our model accommodates that breadth directly.
@@ 7.2 Price Impact
We next characterize the equilibrium price impact matrix $\Lambda$. Its $(i,j)$-element $\Lambda_{i,j}$ is the equilibrium marginal effect of trades in $x_{j}$ on the price of $x_{i}$ (diagonals are within-market impact; off-diagonals are cross impact). The reader can find the notations used in this subsection recalled in Table 1.
@@
Here, $\eta(x_{i}\mid\,\cdot\,)$ is the signal-by-signal payoff profile of claim $x_{i}$. Given order flow $\omega$, the market maker forms a posterior over signals. Throughout the sequel, we write ${\operatorname{\mathbb{Cov}}}(\cdots,\cdots\mid\omega)$ for the (conditional) covariance under that posterior. The outer expectation $\mathbb{E}[\,\cdot\,]$ averages over order flow $\omega$ under its equilibrium distribution. The realized (order-flow-conditional) cross price impact is obtained by dropping the outer expectation from (31):
@@
  • Signal-tilt coefficient $\big(1-\frac{1}{K}\big)$: This is the loading on the realized signal in Step 1 of the informed demand (30) portfolio recipe. When there are many signals ($K$ is large), the realized signal carries less weight relative to the cross-signal average, dampening price impact.
  • Equilibrium aggressiveness $\alpha^{*}$: This is the overall scale of informed demand (30)—Step 3. If the insider trades more aggressively in equilibrium, a given marginal order is more likely to be information-driven, so prices respond one-to-one.
  • Noise adjustment $\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}$: This captures the standard intuition that more noise trading makes order flow less informative. More noise in either market dilutes how much a trade in $x_{j}$ moves beliefs (and thus prices) about $x_{i}$. Indeed, we already saw this in the toy three-state model of Section 2 (see footnote 8).
@@
    is the (posterior) payoff covariance between securities $x_{i}$ and $x_{j}$—and thus captures whether the market maker’s inference from order flow tends to move their prices in the same direction.
@@
    (Negative cross price impact.) If security $x_{i}$ pays off only under signal $s_{k_{i}}$ and $x_{j}$ only under a different signal $s_{k_{j}}\neq s_{k_{i}}$, then their cross impact is negative: $\Lambda_{i,j}<0$. Buy pressure in $x_{i}$ makes the market maker infer that $x_{j}$ is less likely to pay off (and vice versa). More precisely, the sign of $\Lambda_{i,j}$ is determined by the (order-flow-conditional) posterior payoff covariance term in (32), which is negative in this case: